OpenStudy (anonymous):

How can I justify whether these are true or false? det(I+A) = 1 + det(A) and det(A^4) = (det(A))^4

OpenStudy (anonymous):

The determinant is homomorphic. det(AB)=det(A)det(B). To justify the second det(A^4)= det(AAAA)=det(A)det(A)det(A)det(A)=det(A)^4

OpenStudy (anonymous):

The first is not true.

OpenStudy (anonymous):

Just show a single counterexample.

OpenStudy (anonymous):

Do you need to prove that the determinant is homomorphic? I can certainly discuss that.

OpenStudy (anonymous):

I don't think I've even discussed homomorphic

OpenStudy (anonymous):

im interested :)

OpenStudy (anonymous):

but I see what you mean by a counter example I actually have another one that im stuck on if det(A) does not equal 0 then A is always expressible as a product of elementary matrices

OpenStudy (anonymous):

I can't seem to find a counter example this but I can't just say it works

OpenStudy (anonymous):

If det(A) is not zero the matrix is invertible. To invert a matrix you perform a set of elementary row operations to take it to its identity. But you can simply invert those operations and compose them to get back to the original matrix.

OpenStudy (anonymous):

okay that's justifable thanks! I think our entire class is on here today haha

OpenStudy (anonymous):

The point is you can always get from the identity to the invertible matrix.

OpenStudy (anonymous):

With elementary row operations. This clear since you can find the inverse and all elementary row operations are by nature invertible.

OpenStudy (anonymous):

To show det(AB)=det(A)det(B) also involves the idea of being able to decompose an invertible matrix into a composition of elementary operations.

OpenStudy (anonymous):

Each elementary operation corresponds to a change in the determinant.

OpenStudy (anonymous):

you mean like, exchanging two rows changes the sign of the det, multiplying a row by a scalar multiplies the det of a matrix by the same scalar, and multiplying a row by a scalar and adding it to another row doesnt change the det right?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

If either A or B is not invertible than clearly AB is not invertible so det(AB)=0 otherwise you can decompose it into elementary operations and its clear its just the product of all changes to the det.

OpenStudy (anonymous):

gotcha, thank you :)